NO_{2} TRACE MEASUREMENTS BY OPTICAL-FEEDBACK CAVITY-ENHANCED ABSORPTION SPEC-TROSCOPY

\underline{I. VENTRILLARD-COURTILLOT}, Th. DESBOIS, T. FOLDES and D. ROMANINI, Laboratoire de Spectrométrie Physique, CNRS UMR5588, Univ. J. Fourier de Grenoble, St Martin d’Hères, France.

In order to reach the sub-ppb NO_{2} detection level required for environmental applications in remote areas, we develop a spectrometer based on a technique introduced a few years ago, named Optical-Feedback Cavity-Enhanced Absorption Spectroscopy (OF-CEAS) [1]. It allows very sensitive and selective measurements, together with the realization of compact and robust set-ups as was subsequently demonstrated during measurements campaigns in harsh environments [2]. OF-CEAS benefits from the optical feedback to efficiently inject a cw-laser in a V-shaped high finesse cavity (typically 10 000). Cavity-enhanced absorption spectra are acquired on a small spectral region (∼1 cm^{−1}) that enables selective and quantitative measurements at a fast acquisition rate with a detection limit of several 10^{−10} cm^{−1} as reported in this work. Spectra are obtained with high spectral definition (150 MHz highly precisely spaced data points) and are self calibrated by cavity rind-down measurements regularly performed (typically every second). NO_{2} measurements are performed with a commercial extended cavity diode laser around 411 nm, spectral region where intense electronic transitions occur. We will describe the set-up developed for in-situ measurements allowing real time concentration measurements at typically 5 Hz; and then report on the measurements performed with calibrated NO_{2} reference samples to evaluate the linearity of the apparatus. The minimum detectable absorption loss is estimated by considering the standard deviation of the residual of one spectrum. We achieved 2x10^{−10} cm^{−1} for a single spectrum recorded in less than 100 ms at 100 mbar. It leads to a potential detection limit of 3x10^{8} molecules/cm^{3}, corresponding to about 150 pptv at this pressure.

[1] J. Morville, S. Kassi, M. Chenevier, and D. Romanini, Appl. Phys. B, 80, 1027 (2005).

[2] D. Romanini, M. Chenevrier, S. Kassi, M. Schmidt, C. Valant, M. Ramonet, J. Lopez, and H.-J. Jost, Appl. Phys. B, 83, 659 (2006).